OpenStudy (lena772):

The roots of f(x) = 0 are 1 and -2. What are the roots of f(2x)=0? 1 and -2 1/2 and -1 -1/2 and 1 2 and -4 -2 and 4

OpenStudy (lena772):

@ranga

OpenStudy (ranga):

Roots of f(x) are 1 and -2 suggests that f(x) =a(x-1)(x+2) f(2x) = a(2x-1)(2x+2). Its roots are when a(2x-1)(2x+2) = 0 which happens when: x = 1/2 and x = -1.